OpenStudy (anonymous):

Iris has been studying an invasive population of snails. This particular snail has no local predators so the population grows wildly. She has observed that the population follows an exponential rate of growth for fifteen years. Create your own exponential function, f(x), which models the snail population. You will need to identify the principal population of the snails and the rate of growth each year. Explain to Iris how your function shows the principal population and the rate of growth, in complete sentences.

OpenStudy (anonymous):

\[g(x) = 200(1.03)^{2x}=200[(1.03)^2]^x\] so first thing you want to do is square \(1.03\) i.e. compute \[(1.03)(1.03)\]with a calculator that will give you the answer to the first part

OpenStudy (anonymous):

we are looking at part 1 right? the part that says A local snail population grows according to the function g(x) = 200(1.03)2x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent.

OpenStudy (anonymous):

so the first thing you want to do as write it as a function where the exponent is \(x\) instead of \(2x\) clear?

OpenStudy (anonymous):

in other words, instead of \[ g(x) = 200(1.03)^{2x}\] you want \[g(x)=200\times b^x\]for some number \(b\) instead of \(1.03\) your job is to find what \(b\) is

OpenStudy (anonymous):

do you understand that you need to find \(b\)? i am not asking if you know what \(b\) is, only if you know that is what you need to find

OpenStudy (anonymous):

ok now by the laws of exponents, \[(1.03)^{2x}=((1.03)^2)^x\]

OpenStudy (anonymous):

so in order to find the \(b\) what you need to do is find \((1.03)^2=(1.03)(1.03)\) i would use a calculator

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

use a calculator

OpenStudy (anonymous):

close did you use a calculator?

OpenStudy (anonymous):

then make sure to write the answer correctly it is not \(1.069\)

OpenStudy (anonymous):

ok good so now we know \[200(1.03)^{2x}=200(1.0609)^x\]

OpenStudy (anonymous):

that accomplishes job 1, write with an \(x\) in the exponent lol yes, we only did one thing so far... multiply \((1.03)\times (1.03)\)

OpenStudy (anonymous):

i thought we were doing this A local snail population grows according to the function g(x) = 200(1.03)2x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent. Then explain to Iris how the key features of this local snail population compares to the key features of the invasive population.

OpenStudy (anonymous):

we are, the thing on top is just information about iris.

OpenStudy (anonymous):

the first one says to make up your exponential function for a snail poplulation lets do that quickly how many snails would you like to start with?

OpenStudy (anonymous):

ooh, no you did the second part.

OpenStudy (anonymous):

the first part opf the problem is Create your own exponential function, f(x), which models the snail population. You will need to identify the principal population of the snails and the rate of growth each year. Explain to Iris how your function shows the principal population and the rate of growth, in complete sentences.

OpenStudy (anonymous):

uhm..... 5 snails i guess.

OpenStudy (anonymous):

ok fine what percent would you like the population to increase per year?

OpenStudy (anonymous):

5%

OpenStudy (anonymous):

ok good now we have almost answered question one completely Create your own exponential function, f(x), which models the snail population. You will need to identify the principal population of the snails and the rate of growth each year. a) the "principle population is 5" b) the growth rate per year is \(5\%\) c) the function \(f(x)\) will be \(f(x)=5(1.05)^x\)

OpenStudy (anonymous):

in complete sentences, the principle population is 5 which is why out front there is a 5 in \[f(x)=\color{red}5(1.05)^x\]

OpenStudy (anonymous):

the growth rate per year is \(5\%=.05\) to increase a number by \(5\%\) you multiply it by \(100\%+5\%=105\%=1.05\)

OpenStudy (anonymous):

to do it again and again for \(x\) years you multiply by \((1.05)^x\) which is why the function is \[f(x)=5(1.05)^x\] and now we are done with that part completely

OpenStudy (anonymous):

so, for part one for the ansswer i put 100%+5%=105%=1.05 f(x)=5(1.05)x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im asking you

OpenStudy (anonymous):

it asked for "complete sentences" i wrote a couple above, but you can write them in your own words

OpenStudy (anonymous):

hold on im gonna make mmy own sentences and then say then to you and i want you to make sure it sounds right is that ok?

OpenStudy (anonymous):

sure but i gotta run in ten minutes, so be brief

OpenStudy (anonymous):

well since you gatta gho soon lets just move on, because i really want to finish this.

OpenStudy (anonymous):

do we need to redo part 2?

OpenStudy (anonymous):

no that part is good \[200(1.03)^{2x}=200(1.0609)^x\] because \[(1.03)(1.03)=1.0609\]

OpenStudy (anonymous):

so third part.t

OpenStudy (anonymous):

as for Then explain to Iris how the key features of this local snail population compares to the key features of the invasive population. i would say that comparing \[g(x)=200(1.0609)^x\] so \[f(x)=5(1.05)^x\] the first one starts with 200 and grows at a rate of \(6.09\%\) per year, whereas the second one starts with 5 and grows at a rate of \(5\%\) per year

OpenStudy (anonymous):

Iris wants to graph the invasive snail population to show the city council. Justify what the appropriate domain and range would be for the function f(x), what the y-intercept would be, and if the function is increasing or decreasing.

OpenStudy (anonymous):

domain, since \(x\)represents time, we would say that the domain would be \(x>0\) since you can't go backwards in time

OpenStudy (anonymous):

range would be \(y\geq 5\) since we started with 5

OpenStudy (anonymous):

the \(y\)intercept is how many snails you started with you said 5, so the y intercept is 5

OpenStudy (anonymous):

and the function is increasing, because the population is getting bigger, not smaller

OpenStudy (anonymous):

im confused on what i should write for part three though.

OpenStudy (anonymous):

In five years, a garden festival plans on using the park where Iris has been studying the invasive snails. Explain to the garden festival committee how to find the average rate of change for the snail population between years 2 and 5. Describe what this average rate of change represents.

OpenStudy (anonymous):

i take it we are using \[f(x)=5(1.05)^x\]for this one, right?

OpenStudy (anonymous):

wait, im still stuck on three.

OpenStudy (anonymous):

all i have for three is what the domain and range are.

OpenStudy (anonymous):

that is what it asks for

OpenStudy (anonymous):

?

OpenStudy (anonymous):

oh it also asks if it is increasing the population is growing, so it is increasing

OpenStudy (anonymous):

thats all it asks for ?

OpenStudy (anonymous):

alright last part.

OpenStudy (anonymous):

In five years, a garden festival plans on using the park where Iris has been studying the invasive snails. Explain to the garden festival committee how to find the average rate of change for the snail population between years 2 and 5. Describe what this average rate of change represents.

OpenStudy (anonymous):

if we are using \(f(x)=5(1.05)^x\) in year 2 there will be \(5(1.05)^2\)snails and in year 5 there will be \(1.05)^5\) snails compute these numbers first

OpenStudy (anonymous):

then the average is \[5(1.05)^5-5(1.05)^2\]divided by the number of years, which is 3, since \(5-2=3\)

OpenStudy (anonymous):

hey sorry my internet cut off

OpenStudy (anonymous):

uhm what would tyou like me to do my computer cut me off, and im lost now.

OpenStudy (anonymous):

so 5(1.05)2 you want me to solve and 5(1.05)5?

OpenStudy (anonymous):

your messages keep disapearing.

OpenStudy (anonymous):

for 5(1.05)2 the answer i got for that is 27.5625

OpenStudy (anonymous):

@satellite73